OpenStudy (anonymous):

Metal For anyone who can give me the correct answer. write an equation of a line that is perpendicular to the given line and passes through the given points. y - 4 = 5/2 (x + 3); (-7, 8)

OpenStudy (anonymous):

A. Y - 8 = -2/5 (x - 7) B. Y - 8 = 2/5 (x - 7) C. Y - 8 = -2/5 (x + 7) D. Y - 8 = 2/5 (x + 7)

OpenStudy (anonymous):

Please HELP!!!!!

OpenStudy (anonymous):

When two lines are perpendicular, the slope of one is the negative reciprocal of the other.

OpenStudy (anonymous):

So. If the slope of one line is m, the slope of the perpendicular is \[-\frac{1}{m}\]